The vertical height of the point P above the ground is twice that of Q...a particle is projected downwards with a speed of 5m/s from P and at the same time another particle is projected upwards with the same speed from Q....Both particles reach the ground simultaneously,then:a:PQ=b:time of flight of stones=


Grade:10

2 Answers

Jagjot singh
5 years ago
Vq=Vo=V=5m/s (speed of particles)
tq=to=t (time taken to reach the ground)
P=vt+1/2at^2................(i)
Q=-vt+gt^2/2.................(ii)
 
P=2Q..............(iii)
 
Vt+1/2gt^2=2(-vt+gt^2/2)
Vt+gt^2/2=-2vt+gt^2
 
3vt=gt/2
 
t=8v/g
t=8×5/10
t=4sec
Karan Gupta
5 years ago
Let height of point Q from ground = h
Then, height of point P from ground = 2h
 
For point Q, h  =  -5t + (1/2)gt^2.............................(1)
For point P, 2h =  5t + (1/2)gt^2..............................(2)
 
Subtracting (1) from (2), we get: h = 10t
Putting h = 10t in Equation (1), 
10t= -5t + (1/2)*10*t^2
15 = 5t means time of flight T = 3 seconds
and Distance PQ = h = 10t = 30 m